# Impact of interface defects on the band alignment and performance of TiO2/MAPI/Cu2O perovskite solar cells

### v1

@@ IV Conclusions
The performance of PSCs with Cu2O as inorganic HTL in the presence of point defects was investigated using ab initio calculations and SCAPS simulations. In the first step, the band alignment in TiO2/MAPI/Cu2O based PSCs was determined by DFT calculations. Different types of vacancies induce either $n$- or $p$-type doping in the ETL, HTL and MAPI layers, which impact the overall solar cell performance. It was established that in a mostly flat band condition for electrons at the ETL-PRV interface and for holes at PRV-HTL interface, the dominant vacancies are $V_{\rm Ti}$ and $V_{\rm O}$ in the two oxide layers, TiO2 and Cu2O, respectively, while for MAPI $V_{\rm I}$ impacts the interface with the ETL and $V_{\rm Pb}$ influence the interface with the HTL. Therefore, passivation of titanium vacancies in TiO2 and oxygen vacancies in Cu2O should be prioritized in nearly flat-band PSCs. The SCAPS simulations confirm this perspective, by analyzing the PCE dependence on the defect concentration, type and layer localization. To conclude, we performed a comprehensive study on the impact of vacancy defects and provided clues to mitigate potential performance decrease by unintentional defect-induced doping of the active layers.

### v2

@@ IV Conclusions
The performance of PSCs with Cu2O as inorganic HTL in the presence of point defects was investigated using ab initio calculations and SCAPS simulations. In the first step, the band alignment in TiO2/MAPI/Cu2O based PSCs was determined by DFT calculations. Different types of vacancies induce either $n$- or $p$-type doping in the ETL, HTL and MAPI layers, which impact the overall solar cell performance. It was established that in a mostly flat band condition for electrons at the ETL-PRV interface and for holes at PRV-HTL interface, the dominant vacancies are $V_{\rm Ti}$ and $V_{\rm O}$ in the two oxide layers, TiO2 and Cu2O, respectively, while for MAPI $V_{\rm I}$ impacts the interface with the ETL and $V_{\rm Pb}$ influence the interface with the HTL. Therefore, the passivation of titanium vacancies in TiO2 and oxygen vacancies in Cu2O should be prioritized in nearly flat-band PSCs. The SCAPS simulations confirm this perspective, by analyzing the PCE dependence on the defect concentration, type and layer localization. To conclude, we performed a comprehensive study on the impact of vacancy defects and provided clues to mitigate potential performance decrease by unintentional defect-induced doping of the active layers.